OpenStudy (anonymous):

Find the equation of the normal line to the curve f(x) = x^3 - 3x + 1 at the point (2,3) I believe I have to do the derivative which will be the slope correct?

OpenStudy (anonymous):

The derivative I got was 3x-3

OpenStudy (danjs):

derivative is 3x^2 - 3

OpenStudy (anonymous):

oh right I forgot to subtract 1

OpenStudy (danjs):

Using that point, what is the slope of the TANGENT at that point

OpenStudy (anonymous):

is it 9?

OpenStudy (solomonzelman):

~ derivative. ~ plug in the x-coordinate to find the slope. ~ take the perpendicular slope (after step 2) ~ find the equation with the new slope and point (2,3)

OpenStudy (danjs):

m = 3(2)^2 -3 = 9

OpenStudy (danjs):

the slope of a line perpendicular to the TANGENT, will be the slope of the NORMAL at that point

OpenStudy (danjs):

-1/9

OpenStudy (danjs):

with point (2,3) and slope -1/9, form a line

OpenStudy (anonymous):

so one of the equations is y-3 = 9 (x-2) and the normal line is y-3 = -1/9 (x-2) right?

OpenStudy (danjs):

yep

OpenStudy (anonymous):

thank you so much you guys !!!

OpenStudy (danjs):

wait till you start doing those with vectors in 3-space, fun messy as all heck derivatives

OpenStudy (solomonzelman):

\(\large\color{black}{ f(x) = x^3 - 3x + 1 }\) \(\large\color{black}{ f'(x) = 3x^2 - 3 }\) \(\large\color{black}{ f'(2) = 3(2)^2 - 3 =9 }\) \(\large\color{black}{ m_{normal}= -1/9 }\) good!

OpenStudy (danjs):

Tangent, Normal, BiNormal vectors, form a mini coordinate system at any point in 3-space

OpenStudy (solomonzelman):

forgot to put 1/5 spacing here

OpenStudy (solomonzelman):

1.5*